OpenStudy (anonymous):

What is the perimeter of a rectangle with a length of 7 3/4 inches and a width of 4 1/3 inches? A. 24 1/6 inches B. 24 1/3 inches C. 12 1/12 inches D. 11 4/7 inches

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

\(\bf 7\frac{3}{4}+7\frac{3}{4}+4\frac{1}{3}+4\frac{1}{3}=\cfrac{\square ?}{\square ?}\)

OpenStudy (anonymous):

24 1/6

OpenStudy (jdoe0001):

what did you get for the mixed 7 3/4 with only 1 numerator?

OpenStudy (anonymous):

easy you add 3/4+3/4 and that comes to 6/4 and 1/3+1/3=2/3 7+7+4+4=14+8=22 6/4+2/3=22 1/6

OpenStudy (jdoe0001):

well that's actually close to correct 6/4+2/3=2 1/6 thus \(\bf 7\frac{3}{4}+7\frac{3}{4}+4\frac{1}{3}+4\frac{1}{3}=22+2\frac{1}{6}\implies 24\frac{1}{6}\)